OpenStudy (anonymous):

Solve.. 3x^3 - 23x^2 +7x +5

OpenStudy (anonymous):

\[ (1+3 x) \left(5-8 x+x^2\right) \]

OpenStudy (anonymous):

Hi Math, how do we solve this?

hero (hero):

Good job posting the steps Fool

OpenStudy (anonymous):

i mean what is x equal to?

OpenStudy (anonymous):

i think i know the answer now :) thanks!

hero (hero):

cherrydrops, in order to solve it, it has to be an equation. What you have posted is simply an expression.

OpenStudy (anonymous):

are there 3 answers for x? omg the expression has an ' = 0 ' at the right. oops!

hero (hero):

ya, you forgot to include the = 0

OpenStudy (anonymous):

thank you

hero (hero):

FFM, do you mind posting the steps?

OpenStudy (anonymous):

I.. don't think he's there :)

hero (hero):

Yeah, he magically disappeared without posting the steps

OpenStudy (anonymous):

Heheh. yeah :(

hero (hero):

Okay, I have it finally

hero (hero):

Are you ready for the steps?

hero (hero):

3x^3-23x^2+7x+5 3x^3-24x^2+x^2+15x-8x+5 3x^3-24x^2+15x+x^2-8x+5 3x(x^2-8x+5)+1(x^2-8x+5) (x^2-8x+5)(3x+1)

hero (hero):

Tedious, but it works

OpenStudy (anonymous):

wow.. ;) thanks

hero (hero):

now you can do this: (x^2-8x+5)(3x+1) = 0 x^2-8x + 5 = 0 3x+1 = 0 and solve for x

OpenStudy (anonymous):

Aahh i'm lost even before i got to the third step.. ;l how did 3x^3-23x^2+7x+5 become 3x^3-24x^2+x^2+15x-8x+5?

hero (hero):

-24x + x^2 = -23x 15x-8x = 7x

OpenStudy (anonymous):

right... ;